# IndiTag: An Online Media Bias Analysis System Using Fine-Grained Bias Indicators for News Consumers

### v2

@@ 4.3 Evaluation and Validation
The IndiTag system undergoes thorough evaluation to ensure its effectiveness and reliability in analyzing bias in online media content. Evaluation metrics such as precision, recall, and F1-score are utilized to assess the system’s performance, providing insights into its utility for users. Our evaluation of model performance encompasses two key aspects:
@@
2) MicroF1 and MacroF1 for Both Biased and Non-Biased Instances: These metrics provide a comprehensive assessment of the overall prediction capabilities of the models, considering both biased and non-biased instances.
@@
This evaluation framework ensures that the IndiTag system delivers reliable and accurate analysis bias aligned with humans in online media content, thereby empowering users to make informed assessments and interpretations.
4.4 Experimental Results
@@
The comparison results of bias prediction are reported in Table˜1. Across all datasets, the IndiTag generally outperforms other baselines, showcasing its robustness in bias prediction tasks. Notably, IndiTag achieves competitive scores in terms of precision, recall, and F1 scores, indicating its effectiveness in identifying biased instances while maintaining a balance between precision and recall. Moreover, when considering the Micro F1 and Macro F1 scores for both biased and non-biased instances, the IndiTag demonstrates superior performance compared to other baselines. This suggests that IndiTag not only excels in detecting bias but also maintains high prediction capabilities across diverse instances, showcasing its versatility and reliability in bias prediction tasks.
5 User Interface
@@
We present key pages of our IndiTag interface for users in Fig.˜3 for reference. Fig.˜3(a) demonstrates that IndiTag accommodates both single article input and structured document uploading. Fig.˜3(b) and Fig.˜3(c) display the generated descriptors and matched indicators for each descriptor in the input article. The bar beneath each descriptor represents the distribution of probabilities for “left”, “neutral”, and “right” leanings. Additionally, the similarity between each indicator and descriptor is visually represented to aid users in bias analysis. Fig.˜3(d) introduces an additional feature enabling users to input personalized descriptors and articles to analyze the article from the user’s perspective (i.e., descriptor).
6 Conclusion
@@
Despite its strengths, IndiTag also has certain limitations that warrant consideration. One limitation is the reliance on pre-existing bias indicators, which may not encompass the full spectrum of biases present in digital content. Additionally, the effectiveness of IndiTag may vary depending on the quality and diversity of the input datasets used for training and validation. Furthermore, while our system provides valuable insights into bias within digital media, it is not immune to errors or biases inherent in the underlying algorithms and models. Finally, IndiTag’s performance may be influenced by factors such as language nuances and cultural contexts, which may limit its applicability across diverse settings.

### v3

@@ IV-C Evaluation and Validation
The IndiTag system undergoes thorough evaluation to ensure its effectiveness and reliability in analyzing media content bias. Evaluation metrics such as precision, recall, and F1-score are utilized to assess the system’s performance, providing insights into its utility for users. Our evaluation of model performance encompasses two key aspects:
@@
2) MicroF1 and MacroF1 for Both Biased and Non-Biased Instances: These metrics provide a comprehensive assessment of the overall prediction capabilities of the models, considering both biased and non-biased instances. This evaluation framework ensures that the IndiTag system delivers reliable and accurate analysis bias aligned with humans in online media content, thereby empowering users to make informed assessments and interpretations.
IV-D Experimental Results
@@
The comparison results of bias prediction are reported in Table I. Across all datasets, the IndiTag generally outperforms other baselines, showcasing its robustness in bias prediction tasks. Notably, IndiTag achieves competitive scores in terms of precision, recall, and F1 scores, indicating its effectiveness in identifying biased instances while maintaining a balance between precision and recall. Moreover, when considering the Micro F1 and Macro F1 scores for both biased and non-biased instances, the IndiTag demonstrates superior performance compared to other baselines. This suggests that IndiTag not only excels in detecting bias but also maintains high prediction capabilities across diverse instances, showcasing its versatility and reliability in bias prediction tasks.
It is worth noting that in Table I, we compared several models fine-tuned on the FlipBias dataset. These models tend to perform well only on FlipBias itself or on MFC, which shares a similar article-level structure. In contrast, although IndiTag is also developed based on FlipBias, it achieves consistently strong performance across all datasets, demonstrating its high adaptability. Moreover, IndiTag is constructed without a fine-tuning process, allowing it to be easily updated through minor modifications to the indicator database as media bias patterns evolve over time.
IV-E Ablation Study
One of the key improvements from IndiVec to Inditag is the introduction of the strict clustering module. The results of the ablation study are presented in Table II. As shown, applying our indicator framework leads to a significant increase in recall scores, while the addition of the strict clustering module further improves the F1 scores, resulting in a more stable and reliable system.
TABLE II: Ablation study results (in %) on four datasets.
Base Models
FlipBias
BASIL
Pre Rec F1 Pre Rec F1
Full Model 62.9 72.4 67.3 35.2 34.2 34.7
   - $\mathcal{I}$ Clustering 62.7 71.6 66.9 32.2 34.9 33.5
       - $\mathcal{I}$ Indicator 63.9 22.1 32.9 39.3 2.3 4.4
Base Models
BABE
MFC
Pre Rec F1 Pre Rec F1
Full Model 66.4 69.1 67.7 86.4 80.0 83.1
   - $\mathcal{I}$ Clustering 62.9 78.9 70.0 86.9 78.6 82.5
       - $\mathcal{I}$ Indicator 81.9 20.1 32.2 86.5 37.2 52.3
V User Interface
V-A User Interface Design
@@
We present key pages of our IndiTag interface for users in Fig. 3 for reference. Fig. 3(a) demonstrates that IndiTag accommodates both single article input and structured document uploading. Fig. 3(b) and Fig. 3(c) display generated descriptors and matched indicators for each descriptor in the input. The bar beneath each descriptor represents the distribution of probabilities for “left”, “neutral”, and “right” leanings. Additionally, the similarity between each indicator and descriptor is visually represented to aid users in bias analysis. Fig. 3(d) introduces an additional feature enabling users to input personalized descriptors and articles to analyze the article from the user’s perspective (i.e., descriptor).
V-B Case Study
Fig. 3 presents a case illustrating the detailed process of system usage from the perspective of end users. The original news text is as follows:
A Joe Biden presidency could reset ties with top US trade partner Mexico that have suffered since Donald Trump made his first White House bid, tarring Mexican migrants as rapists and gun runners and vowing to keep them out with a border wall.
During the descriptor generation phase, the LLM accurately summarizes this passage with statements such as “Describes Trump’s actions negatively and emphasizes negative impact on US-Mexico relations” and “Focuses on negative aspects of Trump’s actions without presenting counterarguments”. Subsequently, our system matches these descriptors with the bias indicators in the database, such as “Using negative language to describe Donald Trump’s actions and behavior”. Most matched indicators are associated with left-leaning labels, leading the system to assign a left-leaning prediction to the text. Furthermore, the interface visually highlights different leanings using distinct colors, allowing users to easily interpret and verify the analysis results. This visualization provides an intuitive, user-friendly experience for studying media bias.
VI Conclusion
@@
Despite its strengths, IndiTag also has certain limitations that warrant consideration. One limitation is the reliance on pre-existing bias indicators, which may not encompass the full spectrum of biases present in digital content. Additionally, the effectiveness of IndiTag may vary depending on the quality and diversity of the input datasets. Furthermore, while our system provides valuable insights into bias within digital media, it is not immune to errors or biases inherent in the underlying algorithms and models. As large language models continue to evolve, further experiments can be conducted to examine whether the model currently used still outperforms other models. Finally, IndiTag’s performance may be influenced by factors such as language nuances and cultural contexts, which may limit its applicability across diverse settings.
@@
Acknowledge
This work is partially supported by Hong Kong RGC GRF No. 14206324 and CUHK Knowledge Transfer Project Fund No. KPF23GWP20. Thanks for the support of the MoE Key Laboratory of High Confidence Software Technologies (MoE Lab). We are also grateful to the anonymous reviewers for their comments.